HIGHLY PREDISSOCIATIVE LEVELS OF THE CH_{3}S A^{2}A_{1} STATE DETERMINED WITH DEGENERATE FOUR-WAVE MIXING

\underline{YUAN-PERN LEE} and YOSHIYUKI MATSUDA, Department of Chemistry, National Tsing Hua University, Hsinchu, Taiwan 30013.

We have determined highly predissociative levels of the A ^{2}A_{1} state of CH_{3}S in a supersonic jet with degenerate four-wave mixing (DFWM) technique. The highest level observed lies 4265 cm^{−1} above the zero-point-energy level, much greater than corresponding values of 2979 cm^{−1} observed by fluorescence depletion spectroscopy and 1490 cm^{−1} by laser-induced fluorescence. Unlike in fluorescence spectra, relative intensities of lines in DFWM spectra closely reflect their Franck-Condon factors; vibronic assignments are thus more straightforward. A new progression involving excitation of the CH_{3} stretching mode is identified, and several lines in the range 1290-1410 cm^{−1} above the origin may be assigned to the CH_{3} deformation (ν_{5}) and the first overtone of CH_{3} rocking (2ν_{6}) modes. Observed vibrational wave numbers are consistent with theoretical predictions.